OpenStudy (anonymous):

N is the set of real numbers that are factors of 12 a. {1, 2, 3, 4, 6} b. {1, 2, 3, 4, 6, 12} c. {1, 2, 3, 4, 5, 6, 8, 12} d. {1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12}

OpenStudy (perl):

factors of 12 are numbers which you can divide and get a number

OpenStudy (anonymous):

sry its b

OpenStudy (perl):

12/1 = 12 , 12/2 = 6 , 12/12 = 1 , etc so 1, 2, 12 are factors of 12

OpenStudy (perl):

factors of 12 are numbers which you can divide and get a number without a decimal (or remainder )

OpenStudy (anonymous):

so its B?

OpenStudy (anonymous):

I'm confused